OpenStudy (anonymous):

If the density of a certain spherical atomic nucleus is 1.0 × 1014 g cm–3 and its mass is 2.0 × 10–23 g, what is its radius in cm? I need to understand how to do it not only the answer please :)

OpenStudy (jfraser):

knowing the density and mass allows you to find the volume of the nucleus, which we assume is spherical in shape:\[density = \frac{mass}{volume}\]so\[Volume = \frac{mass}{density}\]The volume of a sphere is always given by:\[Volume = \frac{4}{3} *\pi * r^3\]Using the volume you solve for earlier, solve for the radius second.